# Comparing Hallucination Detection Metrics for Multilingual Generation

### v1

@@ Abstract
While many automatic hallucination detection techniques have been proposed for English texts, their effectiveness in multilingual contexts remains unexplored. This paper aims to bridge the gap in understanding how these hallucination detection metrics perform on non-English languages. We evaluate the efficacy of various detection metrics, including lexical metrics like ROUGE and Named Entity Overlap and Natural Language Inference (NLI)-based metrics, at detecting hallucinations in biographical summaries in many languages; we also evaluate how correlated these different metrics are to gauge whether they measure the same phenomena. Our empirical analysis reveals that while lexical metrics show limited effectiveness, NLI-based metrics perform well in high-resource languages at the sentence level. In contrast, NLI-based metrics often fail to detect atomic fact hallucinations. Our findings highlight existing gaps in multilingual hallucination detection and motivate future research to develop more robust detection methods for LLM hallucination in other languages.
@@ 1 Introduction
Large Language Models (LLMs) have brought about remarkable advances in text generation. However, they are still prone to factuality hallucination - the generated text conflicts with established world knowledge Huang et al. (2023); Zhang et al. (2023). Despite considerable research efforts towards understanding and detecting hallucinations, the focus has predominantly been on English texts Huang et al. (2023); Zhang et al. (2023); Ji et al. (2023). This emphasis has resulted in a significant knowledge gap regarding hallucinations in multilingual contexts, and it is currently unclear whether the methods developed for detecting and addressing hallucinations in English are effective or even applicable in multilingual settings.
@@
In this paper, we address the highlighted issues by evaluating the effectiveness of various metrics (initially proposed for English generation) within a multilingual context. Our approach involves a comparative analysis of traditional lexical metrics, such as ROUGE Lin (2004) and Named Entity Overlap, alongside Natural Language Inference (NLI) metrics. Additionally, we compare reference-based metrics with pairwise metrics that are based on the consistency among generated samples; we also present a correlation study between these automated metrics and human factuality verification. These evaluations shed light on the efficacy of these metrics in non-English languages and aim to establish a more robust framework for hallucination detection in multilingual language models.
We empirically evaluate different hallucination detection techniques in the multilingual context and find that:
  • Lexical overlap metrics (e.g., ROUGE, named entity overlap) do not correlate with NLI-based metrics in detecting factual hallucinations in both reference-free and reference-based settings.
  • Pairwise NLI-based metrics strongly correlate with reference-based metrics in high-resource languages, but this significantly diminishes in low-resource languages.
  • Automatic NLI-based metrics effectively identify sentence-level hallucinations in high-resource languages when compared to human evaluations. However, their performance diminishes when assessing simpler atomic facts.
@@
These findings indicate several open problems for multilingual hallucination detection. While traditional lexical overlap methods and pairwise comparisons of multiple generated texts are more accessible approaches for low-resource languages, they are often inadequate for hallucination detection. This highlights that the effectiveness of hallucination detection is closely tied to the availability and quality of language resources, following the trend observed in English that detection accuracy depends on natural language understanding abilities Manakul et al. (2023); Min et al. (2023). Our work points to a substantial gap in hallucination detection in multilingual contexts, necessitating focused future research to bridge this divide.
2 Analyzing Methods for Multilingual Hallucination Detection
Hallucination refers to machine-generated content that is not faithful to a reference. This reference could be source text, preceding generated context, or world knowledge Ji et al. (2023); Zhang et al. (2023). Hallucinations can be further classified into two types: verifiable hallucinations, which happen when the generated content directly contradicts the reference, and unverifiable hallucinations, meanwhile, refer to generated content that cannot be verified with the source material.
@@ 2 Analyzing Methods for Multilingual Hallucination Detection
In this paper, we measure the efficacy of different metrics on detecting multilingual hallucinations. We focus on biography generation, a domain that is particularly sensitive to factual accuracy and coherence Min et al. (2023); Dhuliawala et al. (2023). We test a suite of automatic metrics, each of which caters to a different aspect of factual generation: ROUGE Lin (2004), named entity overlap, and Natural Language Inference (NLI)-based methods.
@@ 2.1 Multilingual Biography Generation
Table 1: Generation quality statistics for BLOOMZ-mt. The languages in bold are the most frequent in the ROOTS pretraining corpus Laurençon et al. (2022), and the underlined languages are covered in the xP3mt fine-tuning dataset Muennighoff et al. (2023). ”FLang.” refers to the most frequently generated language for each prompt language.
@@
The length of the generated texts varies notably across languages. High-resource languages like English and French average 78.3 and 115.8 tokens per response, respectively. However, mid-resource languages such as Thai tend to generate much shorter biographies and incomplete sentences, and low-resource languages fare even worse (for instance, Ukrainian averages just 5.7 tokens and 0.40 sentences), demonstrating the significant gap in generation abilities across languages.
We assess the accuracy of the generated languages through three metrics: the percentage of valid generations that is detectable for the langdetect package (Valid %) [^1] Some generations are incomplete, or their language is undetectable., the most frequently generated language for a given target language (Flang), and the accuracy of generated language out of the valid generations (Acc.). For high-resource languages like English, Chinese, Spanish, and French, the models generally generate text in the correct language; however, for the languages highlighted with an underwave the model generates in the wrong language the majority of the time. Often, this is due to the model generating in a closely related high-resource language. For languages such as Italian and Bulgarian, many inaccurate generations are in English. Similarly, Japanese generations often switch to Chinese when mistakes occur. Languages with more distinctive linguistic features —such as Thai’s unique script—facilitate more accurate model generations.
2.2 Automatic Methods for Multilingual Factuality Detection
Following the language verification and ruling out examples where the generations are in the wrong languages, the next phase involves the detection of hallucinations in a multilingual environment by assessing the consistency between a target generation and either a reference text or its other generations. This section considers various automatic metrics for detecting hallucinations in long-form generations.
@@ 2.2 Automatic Methods for Multilingual Factuality Detection
The ROUGE metric is employed to assess the token-level similarity between texts. We consider the ROUGE 1 (R1), 2 (R2), L (RL), and Lsum (RLsum) scores of the generated text against the reference.
@@
Pairwise We generate $k$ samples for each biography. In this setting, we compare $\hat{t}$ against the other generated samples for the same person and calculate the average score across all generations.
@@ 3 Experiment Setup
3.1 Dataset
Our curated dataset encompasses 19 languages: English, Spanish, Russian, Indonesian, Vietnamese, Persian, Ukrainian, Swedish, Thai, Japanese, German, Romanian, Hungarian, Bulgarian, French, Finnish, Korean, Italian, and Chinese. Using WikiData, we extract 500 human names that are covered by all of these languages on Wikipedia, based on diverse page view counts from 2022-01-01 to 2023-01-01. For our reference text, we use the Wikipedia API to obtain the full-page content. We detect instances where the LLMs generate text in an incorrect language with langdetect. [^2] APIs: https://query.wikidata.org/, https://pypi.org/project/wikipedia/, and https://pypi.org/project/langdetect/, respectively
3.2 Models and Prompting Details
In our experiments, we deploy the pretrained multilingual BLOOM model as well as the BLOOMZ-mt model, which is fine-tuned with machine-translated prompts Workshop (2023). We use nucleus decoding Holtzman et al. (2020) with $top\_p=0.9$ and generate five responses per prompt. For each evaluation language, we generate a prompt template with Google Translate. The template in English is ”Tell me a biography of ¡Name¿.”; all templates are in Appendix (Figure 2).
3.3 Detection Metric Details
@@
# examples in correct language ¿ 1,000
@@
# examples in correct language ¡ 1,000
@@
# examples in correct language ¡ 100
@@
Table 2: The correlation between the reference-based NLI result and the pairwise NLI result across different languages. The languages with underline are covered in the XNLI finetuing dataset. The numbers in yellow have the p-values larger than 0.05.
4 Results
In this section, we compare evaluation metrics to estimate hallucinations in our generated biographical corpus (Section 4.1). We then perform a correlation study to test whether the proposed metrics agree when hallucination occurs (Section 4.2). We find that traditional, reference-based lexical metrics (i.e., ROGUE and named entity overlap) are likely insufficient to capture hallucinations, as they rarely correlate with stronger NLI-based approaches. We also observe that pairwise metrics detect hallucinations less often in lower-resource languages when compared to reference-based metrics.
4.1 Comparison of Automatic Hallucination Metrics
Table 3: Results of different reference-based metrics for the BLOOMZ-mt model. ”-” indicates the language is not covered by the Spacy NER tool. All of the ROUGE and Named Entity Overlap (NEO) results are in percentage (%).
Language R1-F1 R1-P R1-R R2-F1 R2-P R2-R NEO-F1 NEO-P NEO-R $\mathrm{DIFF}$ $\mathrm{UNV}$
High-Resource Languages
English 1.83 87.58 0.94 0.87 47.38 0.44 4.27 53.41 2.26 -0.60 0.19
Chinese 6.43 57.34 3.76 2.07 23.22 1.17 4.69 35.27 2.79 -0.62 0.21
Spanish 2.77 85.86 1.47 1.35 49.10 0.71 3.28 48.48 1.76 -0.51 0.18
French 2.18 87.78 1.13 1.06 51.41 0.55 4.35 57.41 2.31 -0.54 0.16
Vietnamese 6.82 92.92 4.22 4.10 73.28 2.43 - - - -0.49 0.15
Indonesian 7.51 68.51 4.87 2.36 26.39 1.53 - - - -0.45 0.22
Middle-Resource Languages
German 0.38 71.34 0.19 0.11 35.22 0.05 0.83 36.06 0.42 -0.65 0.15
Italian 0.50 69.13 0.25 0.13 29.82 0.07 1.00 30.26 0.52 -0.58 0.17
Japanese 0.73 14.62 0.40 0.05 1.50 0.03 0.47 15.52 0.25 -0.72 0.21
Bulgarian 0.16 4.92 0.09 0.01 0.37 0.01 - - - -0.61 0.19
Romanian 1.02 69.75 0.53 0.42 45.79 0.22 0.39 17.47 0.20 -0.29 0.24
Swedish 0.66 86.37 0.33 0.32 77.23 0.16 1.28 45.24 0.66 -0.40 0.63
Low-Resource Languages
Thai 0.04 1.14 0.02 0.00 0.00 0.00 - - - -0.56 0.38
Russian 0.09 4.69 0.05 0.01 0.28 0.00 0.48 11.28 0.25 -0.58 0.47
Ukrainian 0.04 1.53 0.02 0.00 0.00 0.00 0.70 20.64 0.36 -0.53 0.66
Persian 0.00 0.00 0.00 0.00 0.00 0.00 - - - -0.50 0.92
Finnish 0.89 37.70 0.46 0.20 10.03 0.10 0.58 23.71 0.30 -0.59 0.91
Korean 0.18 6.58 0.09 0.01 0.88 0.00 0.24 8.48 0.12 -0.53 0.94
Hungarian 0.74 64.74 0.37 0.16 23.23 0.08 - - - -0.53 0.97
@@ 4.1 Comparison of Automatic Hallucination Metrics
We first consider how different automatic methods for detecting hallucination perform across languages on the generated biographical data from the BLOOMZ-mt model (Table 3). High-resource languages, such as English, Chinese, Spanish, French, Vietnamese, and Indonesian, exhibit higher recall scores, which suggests that the text generated in these languages has better coverage of the corresponding Wikipedia reference content. This contrasts sharply with lower-resource languages, which demonstrate significantly diminished recall. Interestingly, languages that frequently produce incorrect language outputs (e.g., German and Italian) or often yield empty or incomplete generations (e.g., Swedish and Hungarian) still have relatively high precision scores. While these generations seem to contain few explicit hallucinations, they also exclude many facts from the reference, as indicated by their correspondingly low recall scores.
@@
When considering the $\mathrm{DIFF}$ metrics, all languages under scrutiny resulted in negative scores, which was consistent even among the higher-resource languages like English and Chinese. This indicates a tendency towards contradictions in the generated text with their respective reference texts. For the $\mathrm{UNV}$ scores, higher and middle-resource languages (ranging from English to Romanian in the table 3) cluster within a similar scope of 0.15 to 0.25. In contrast, low-resource languages that often produce empty or incomplete generations, such as Ukrainian, Persian, Finnish, and Korean, show much higher $\mathrm{UNV}$ scores. This implies that the $\mathrm{UNV}$ metric is sensitive to incomplete text generations and missing information and may indicate the model’s generation errors beyond hallucination. As for the reference-free pairwise metrics, we observe similar trends across different languages (Table 9).

### v2

# Comparing Hallucination Detection Methods for Multilingual Generation

@@ Abstract
While many hallucination detection techniques have been evaluated on English text, their effectiveness in multilingual contexts remains unknown. This paper assesses how well various factual hallucination detection metrics (lexical metrics like ROUGE and Named Entity Overlap, and Natural Language Inference (NLI)-based metrics) identify hallucinations in generated biographical summaries across languages. We compare how well automatic metrics correlate to each other and whether they agree with human judgments of factuality. Our analysis reveals that while the lexical metrics are ineffective, NLI-based metrics perform well, correlating with human annotations in many settings and often outperforming supervised models. However, NLI metrics are still limited, as they do not detect single-fact hallucinations well and fail for lower-resource languages. Therefore, our findings highlight the gaps in exisiting hallucination detection methods for non-English languages and motivate future research to develop more robust multilingual detection methods for LLM hallucinations. [^1] The code and annotated dataset will be released upon publication.
@@ 1 Introduction
Large Language Models (LLMs) have made remarkable advances in text generation. However, they are still prone to hallucinating facts, or generating text that conflicts with established world knowledge Huang et al. (2023); Zhang et al. (2023). While there has been considerable research towards detecting hallucinations in English Huang et al. (2023); Zhang et al. (2023); Ji et al. (2023), much less focus has been given to multilingual hallucinations. Therefore, it is currently unclear whether the methods developed for detecting and addressing hallucinations in English are effective or even applicable in multilingual settings.
@@
This paper evaluates the effectiveness of various automatic metrics, initially proposed for English factual hallucination detection, within a multilingual context. We focus on automatic metrics requiring minimal in-language resources to perform hallucination detection, as this makes them most readily applicable to new languages; these metrics include traditional lexical metrics, such as ROUGE Lin (2004) and Named Entity Overlap, as well as Natural Language Inference (NLI) metrics. We also consider the differences between reference-based metrics and pairwise metrics based on the consistency among generated samples. To evaluate these metrics, we present correlation studies comparing these automated metrics directly, against supervised hallucination detection methods, and with human judgments of generation factuality.
We empirically evaluate these hallucination detection techniques in the multilingual context with a new dataset of parallel biographical generations (Section 2.1). Our experiments find that: (1) lexical overlap metrics do not agree with NLI metrics or human judgments when detecting hallucinations in reference or pairwise settings; (2) while pairwise NLI metrics strongly correlate with reference-based ones in high-resource languages, this significantly diminishes in low-resource settings; (3) automatic NLI metrics effectively detect sentence-level hallucinations in high-resource languages when compared to human evaluations, but not when assessing atomic facts; and (4) NLI metrics outperform supervised approaches at detecting hallucinations that can be verified or refuted by the reference text, but not on unverifiable errors.
@@
Overall, while lexical overlap methods and pairwise comparisons of generated texts are more accessible for evaluating low-resource languages, they are often inadequate at hallucination detection. Additionally, while NLI-based metrics can detect factual hallucinations — and even outperform models trained on hallucination detection in some cases — these metrics perform best on high-resource languages. This highlights that multilingual hallucination detection performance is closely tied to the availability and quality of language resources, mirroring the trend observed in English that detection depends on natural language understanding abilities Manakul et al. (2023); Min et al. (2023). This points to a substantial gap in hallucination detection in multilingual and low-resource contexts and the need for future work bridging this divide.
2 Multilingual Hallucination Detection
@@
We measure the efficacy of different automatic metrics on detecting multilingual hallucinations. We focus on biography generation, a domain that is particularly sensitive to factual accuracy and coherence Min et al. (2023); Dhuliawala et al. (2023). We test a suite of automatic metrics, each of which caters to a different aspect of factual generation: ROUGE Lin (2004), named entity overlap, and Natural Language Inference (NLI)-based methods.
@@ 2.1 Multilingual Biography Generation
The length of the generated texts varies notably across languages. While high-resource languages like English and French generate longer outputs, mid-resource languages such as Thai tend to generate much shorter biographies and incomplete sentences. Low-resource languages fare even worse (for instance, Ukrainian averages just 5.7 tokens and 0.40 sentences), demonstrating the significant gap in generation abilities across languages.
We assess the accuracy of the generated languages through three metrics: the percentage of valid generations that is detectable for the langdetect package (Valid %), the most frequently generated language for a given target language (Flang), and the accuracy of generated language out of the valid generations (Acc.). For high-resource languages like English, Chinese, Spanish, and French, the models generally generate text in the correct language; however, for the languages highlighted with an underwave the model generates in the wrong language the majority of the time. Often, this is due to the model generating in a closely related high-resource language. For languages such as Italian and Bulgarian, many inaccurate generations are in English. Similarly, Japanese generations often switch to Chinese when mistakes occur. Languages with more distinctive linguistic features —such as Thai’s unique script—facilitate more accurate model generations.
@@
Table 1: Quality statistics for BLOOMZ-mt generations. Languages that occur in the ROOTS pretraining corpus are in bold Laurençon et al. (2022), and underlined languages are in the xP3mt fine-tuning dataset Muennighoff et al. (2023). "FLang." refers to the most frequently generated language for each prompt language.
2.2 Automatic Metrics
After quality verification of generated samples and filtering examples where the output is in an incorrect language, we compare the efficacy of different hallucination detection metrics on the remaining generations. We consider automatic metrics for detecting hallucinations in long-form generations that work by assessing the consistency between a target generation and either a reference text or its other generations. Specifically, we focus on metrics that do not require supervised hallucination data: many languages do not have datasets available for this task, which makes these supervised methods infeasible for those settings. [^2] However, for completeness we include two recent supervised methods for multilingual hallucination detection in §5.
@@ 2.2 Automatic Metrics
The ROUGE metric is employed to assess the token-level similarity between texts. We consider the generated text’s ROUGE 1 (R1) and L (RL) scores against the reference.
@@
Pairwise We generate $k$ samples for each biography. In this setting, we compare $\hat{t}$ against the other generated samples for the same person and calculate the average score across all generations. Experimental details for calculating these metrics are given in the Appendix.
@@ 3 Experiment Setup
Dataset
Our curated dataset encompasses 19 languages: English, Spanish, Russian, Indonesian, Vietnamese, Persian, Ukrainian, Swedish, Thai, Japanese, German, Romanian, Hungarian, Bulgarian, French, Finnish, Korean, Italian, and Chinese. Using WikiData, we extract the names of 500 people who are covered by all of these languages on Wikipedia, based on diverse page view counts from 2022-01-01 to 2023-01-01. For our reference text, we use the Wikipedia API to obtain the full-page content. We detect instances where the LLMs generate text in an incorrect language with langdetect, which covers all 19 languages in our experiment. [^3] APIs: https://query.wikidata.org/, https://pypi.org/project/wikipedia/, and https://pypi.org/project/langdetect/, respectively.
Models and Prompting
We generate text samples with the BLOOMZ-mt model, which is fine-tuned with machine-translated prompts Workshop (2023); at the time of our experiments, BLOOMZ-mt is the largest open-source, multilingual LM. We use nucleus decoding Holtzman et al. (2020) with $top\_p=0.9$, which is a common and realistic configuration used in other works in LLM hallucination Liu et al. (2023), and generate five responses per prompt to evaluate the pairwise, intrinsic metrics. For each evaluation language, we generate a prompt template with Google Translate. The template in English is "Tell me a biography of <Name>."; the templates translated into other languages are in Appendix (Figure 2).
Language R1-F1 R1-P R1-R RL-F1 RL-P RL-R N-F1 N-P N-R $\mathrm{DIFF}$ $\mathrm{UNV}$ $\mathrm{ENT}$
High-Resource Languages
English 1.83 87.58 0.94 1.40 72.41 0.72 4.27 53.41 2.26 -0.60 0.19 0.16
Chinese 6.43 57.34 3.76 5.59 51.73 3.26 4.69 35.27 2.79 -0.62 0.21 0.16
Spanish 2.77 85.86 1.47 2.19 72.39 1.16 3.28 48.48 1.76 -0.51 0.18 0.28
French 2.18 87.78 1.13 1.67 73.51 0.87 4.35 57.41 2.31 -0.54 0.16 0.25
Vietnamese 6.82 92.92 4.22 5.34 85.87 3.21 - - - -0.49 0.15 0.34
Indonesian 7.51 68.51 4.87 5.44 55.28 3.52 - - - -0.45 0.22 0.32
Middle-Resource Languages
German 0.38 71.34 0.19 0.31 67.60 0.16 0.83 36.06 0.42 -0.65 0.15 0.50
Italian 0.50 69.13 0.25 0.42 63.77 0.21 1.00 30.26 0.52 -0.58 0.17 0.42
Japanese 0.73 14.62 0.40 0.64 13.53 0.35 0.47 15.52 0.25 -0.72 0.21 0.26
Bulgarian 0.16 4.92 0.09 0.15 4.91 0.08 - - - -0.61 0.19 0.50
Romanian 1.02 69.75 0.53 1.00 69.08 0.52 0.39 17.47 0.20 -0.29 0.24 0.76
Swedish 0.66 86.37 0.33 0.64 85.87 0.33 1.28 45.24 0.66 -0.40 0.63 0.79
Low-Resource Languages
Thai 0.04 1.14 0.02 0.04 1.14 0.02 - - - -0.56 0.38 0.32
Russian 0.09 4.69 0.05 0.09 4.62 0.05 0.48 11.28 0.25 -0.58 0.47 0.40
Ukrainian 0.04 1.53 0.02 0.03 1.52 0.02 0.70 20.64 0.36 -0.53 0.66 0.51
Persian 0.00 0.00 0.00 0.00 0.00 0.00 - - - -0.50 0.92 0.38
Finnish 0.89 37.70 0.46 0.80 35.61 0.41 0.58 23.71 0.30 -0.59 0.91 0.33
Korean 0.18 6.58 0.09 0.18 6.57 0.09 0.24 8.48 0.12 -0.53 0.94 0.25
Hungarian 0.74 64.74 0.37 62.56 23.23 0.36 - - - -0.53 0.97 0.51
Table 2: Results of different reference-based metrics for the BLOOMZ-mt model. "-" indicates the language is not covered by the Spacy NER tool. All of the ROUGE and Named Entity Overlap (N) results are in percentage (%).
4 Multilingual Hallucination Metrics
This section compares how different automatic metrics estimate hallucinations in our generated biographical corpus (Section 4.1). We then perform a correlation study to test whether these metrics agree when hallucination occurs (Section 4.2).
4.1 Automatic Metrics
We first consider how different referenced-based automatic methods for detecting hallucination perform across languages on the generated biographical data from the BLOOMZ-mt model (Table 2). [^4] We observe similar trends on pairwise metrics (Table 10). We find that, unsurprisingly, these measures indicate increases in hallucination on middle- and low-resource languages (e.g., lower overlap with the reference, higher $\mathrm{UNV}$ scores). However, the NLI-based $\mathrm{DIFF}$ scores remain relatively stable regardless of language resourcefulness.
@@ 4.1 Automatic Metrics
# examples in correct language > 1,000
@@
# examples in correct language < 1,000
@@
# examples in correct language < 100
@@
Table 3: The correlation between the reference-based NLI result and the pairwise NLI result across different languages. The languages with underline are covered in the XNLI finetuing dataset. The numbers in gray have the p-values larger than 0.05.
Lexical Overlap Metrics
We also note some specific trends within this metric type. For example, high-resource languages (English, Chinese, Spanish, French, Vietnamese, and Indonesian) exhibit particularly high recall scores, suggesting that the text generated in these languages has better coverage of the corresponding Wikipedia reference content. In contrast, lower-resource languages demonstrate significantly diminished recall.
@@
Interestingly, languages where BLOOMZ-mt frequently produces incorrect language outputs (e.g., German and Italian) or empty or incomplete generations (e.g., Swedish and Hungarian) maintain relatively high precision scores in the higher-quality outputs we evaluate. While these generations seem to contain few explicit hallucinations, they often exclude many facts from the reference, as indicated by their correspondingly low recall scores.
NLI-based Metrics
All languages we consider obtain negative $\mathrm{DIFF}$ scores, including higher-resource languages like English and Chinese. This indicates a tendency towards contradictions in the generated text with their respective reference texts — as measured by the NLI classifier.
@@
For the $\mathrm{UNV}$ scores, higher and middle-resource languages (ranging from English to Romanian in the table 2) fall within a similar range of 0.15 to 0.25. In contrast, low-resource languages that often produce empty or incomplete generations, such as Ukrainian, Persian, Finnish, and Korean, obtain much higher $\mathrm{UNV}$ scores. This implies that the $\mathrm{UNV}$ metric is sensitive to incomplete text generations and missing information and may indicate the model’s generation errors beyond hallucination.